### v2

# Strong parity-violation effects induced by large-amplitude motions: A quantum-dynamics study of substituted chiral methanols

Abstract
An enhanced mechanism is proposed for the large-amplitude-motion-induced parity-violating frequency by integrating the exact quantum dynamics method with the relativistic electronic structure theory. The torsional wavefunctions and PV frequency shifts are obtained by the exact quantum dynamics method. The potential energy curve and PV energy along the torsional coordinates are calculated using the extended atomic mean-field two-component Hamiltonian. The predicted PV frequency shift for the torsional transition of CFClBrOH is approximately 100 times larger than that of the conventional C-F stretching mode of CHFClBr. The maximum PV frequency shift (3.1 Hz) is obtained in the CHBrIOH molecule.
I Introduction
What will humans discover after improving the precision of molecular spectra? The advancement of cooling techniques for polyatomic molecules [1, 2, 3, 4] can transcend the precision of traditional spectroscopies. Although the rotation-vibrational spectra are mainly generated by electromagnetic interactions, the parity-violating (PV) interaction may split them. The PV interaction was first observed in nuclear beta decay [5], using the optical-rotation technique [6, 7] as well as in atomic forbidden transition due to a nuclear anapole moment [8]. Various molecules have been investigated as will be discussed. However, only the upper limit of the PV frequency shift has been reported using a methane-like chiral molecule, CHFClBr [9, 10]. Although the PV energy difference between the enantiomer () can appear in the framework of the standard model [11, 12, 13], the molecular PV shift may be a probe of cold dark matter [14, 15].
[…]
Here, the PV shift in the torsional motions of the CXYZ unit in CXYZOH (X, Y, Z = H, F, Cl, Br, I) molecules was investigated. Strong PV enhancements are observed due to the localization of the torsional wavefunction to the geometries exhibiting opposite signs. The results obtained for CHBrIOH, CFBrIOH, CFClBrOH, and CClBrIOH molecules are reported as examples, and those for the other molecular species are reported in the Supplementary Material (SM).
II Theory
Figure 1 shows the flow diagram for obtaining the PV frequency of the torsional states.
- 1.
2.
The total energies (the potential energy curve, PEC) and were calculated using optimized geometries at ; solid black line in Figure 3).
3.
The minimum energy on the PEC was obtained using the simplex code [37], and the PEC was shifted so that its bottom could become zero .
4.
The Cartesian coordinates optimized in Step 1 with respect to were transformed into the relaxed reference internal coordinates. First, we extracted and from the Cartesian coordinate defined in Eq. 4 and obtain the torsional angle (; Eq. 20). Further, (, which covers ) is obtained from , and these discrete coordinates are interpolated to obtain . Additionally, the horizontal axis of the potential energy is also transformed (, while is interpolated.
- 5.
The following definition was used to obtain the body-fixed Cartesian coordinate (Figure 2), following Ref. 37
where the atomic number increases from atoms X-Z. The frame for the internal rotation of the CXYZ unit was also defined based on the body-fixed Cartesian coordinate definition (Eq. 4). Furthermore, the center of mass was fixed at the origin of the body-fixed coordinate system, which offers an additional constant shift vector to each vector in Eq. 4. These coordinate definitions were used in the vibrational calculations to prevent the rotation of the frame.
The electronic structure calculations, including geometry optimization, PEC, and calculations, were performed on structures parametrized by H-C-O-Z dihedral angle, , whereas the torsional motion in the quantum dynamics calculation was described by , defined as a linear combination of and
This definition facilitated the identification of as the single LAM in the system, with the small-amplitude motions (SAMs). This extraction technique for was initially proposed in Refs. [38, 39]. For subsequent convenience, we define the following compact notation for the internal coordinates,
where collects the SAMs. The mathematically allowed coordinate ranges are defined as , and the other coordinates are inactive in this study.
II.1 Vibrational Hamiltonian
The Hamiltonian used to describe the quantum vibrational motion can be expressed in the Podolsky form [40]:
where is the number of the vibrational degree of freedom in the molecule, is the momentum operator with respect to the internal coordinates, , , and is the potential energy depending on the coordinates of the nuclei. In our vibrational calculations, we considered only the torsional motion (i.e., ). In the GENIUSH program [41, 42], the mass-weighted metric tensor, , was calculated from the -vectors, as follows:
with the vibrational -vector:
and the rotational -vector:
where is the unit vector in the body-fixed Cartesian coordinates. In the actual calculations, the -vectors are obtained at the discrete-variable representation (DVR) grid points (), i.e., the matrix elements are constructed based on rather than in eqs 24 and 25. The reference internal coordinates for obtaining are relaxed at each , where is the geometry that minimizes at a given , . Employing the relaxed coordinates indicates that the moment of inertia in this study depends on . The wavefunctions and energies associated with Eq. 22 are obtained in a variational manner.
II.2 Parity-violating Hamiltonian
The PV effective Hamiltonians arise from the coupling between the vector [] and axial-vector [] terms of the Fermion currents [43, 44, 11, 20]. Employing indices of an electron (e) and a nucleon [proton (p) and nucleon (n)], and contribute to the nuclear spin-dependent term and independent term, respectively. The former could appear in nuclear magnetic resonance (NMR) [45, 46, 47, 48] owing to the contribution from the nuclear spin (. The latter is our target, where the effective Hamiltonian in the four-component framework is shown as [43, 44, 11, 20]
Here, is the Fermi coupling constant, is the weak charge of the nucleus , with its neutron number and nuclear charge . The Weinberg mixing angle [49, 50] was also employed. is the Dirac matrix that results in the coupling between the large and small components of the wavefunction. Additionally, is the charge density of nucleus for which the Gaussian-type model [51] is employed in this study. As the PV interaction is mediated by the heavy boson, the contact interaction between and an electron offers a good approximation. Although Eq. 26 is a four-component expression, it is numerically transformed into a two-component framework (picture-change [52]) in this study. As the sign of the becomes opposite after parity inversion, the energy difference due to the PV interaction between the - and -configuration can be expressed by
where is the target configuration of the molecule ( or ). is the PV energy obtained from the expectation value with respect to the electronic wavefunction, . The PV shift at the vibrational state, , is given, as follows:
where expresses the value of at the nuclear configuration, . We obtained the expected value with respect to the DVR vibrational wavefunction. This indicates that the error due to the truncation of the potential (e.g., the harmonic oscillator and anharmonic correction), which is traditionally employed in this field, is not observed in this manner.
III Computational details
The vibrational Schrödinger equation in a reduced-dimension form was solved using the GENIUSH program [41, 42], which was recently reviewed in Ref. 53. The derivative in Eq. 24 was obtained numerically (two-point central derivative) using the step size bohr. Furthermore, 37 Fourier-DVR grid points, which is a sufficient number of grid points as presented in Table S1 of the SM, were used to solve the vibrational Schrödinger equation. The TENSINT program [54, 55] was used to obtain at each torsional energy level.
Geometry optimization was carried out using the Gaussian16 software package [56]. For the Br and I atoms, we employed the pseudopotentials ECP10MDF and ECP28MDF, respectively, and the cc-pVTZ-PP basis sets [57, 58]. For the H, C, O, F, and Cl atoms, we employed the Def2TZVPP basis sets [59]. The PBE1PBE functional [60, 61] was used for the density functional theory (DFT) functional. Dispersion effects were considered within the semi-empirical D3 van der Waals corrections [62].
To calculate the total electronic energy and , we employed a development version (git hash: 3970bcf) of the DIRAC program package [63, 64]. We constructed a two-component relativistic Hamiltonian using the exact two-component (X2C) Hamiltonian [65] for the one-electron operators and the extended atomic mean-field (eAMF-X2C) approach [66] to consider the picture-change effect of the two-electron Coulomb and Gaunt terms. The PBE0 functionals [60, 61] were employed for the electronic structure calculations, and the dyall.3zp basis sets [67, 68, 69] were employed for all the elements in the molecules in the uncontracted form. Several tight exponents, which were optimized using the simplex method [70], were also added (Table S2 of the SM). The Gauss-type finite nuclear-charge model [51] was employed for the relativistic calculations.
IV Results and discussion
Figure 3 shows the criticality of the geometry-relaxation effect in the generation of the 1D PEC and curve. The fixed-geometry model cannot reproduce the results of the relaxed structure, as some wells in the PEC do not appear in the fixed-geometry model. Although the situation is better in , the fixed-geometry model overestimates the PV sensitivities in many regions. From a different perspective, geometries deviating from the optimized torsional path can display larger PV sensitivities, and it encourages full-dimensional quantum dynamics calculations in future studies. The relaxed structure model was used for the following discussion. Although the shape of the curve appeared complicated, two sinusoidal curves (a local minimum and a local maximum at 0∘-180∘ and 180∘-360∘, respectively) can be found for some target molecules (see Figures 4 and S1-S6 in the SM).
The enhancement of the PV frequency shift induced by the torsional motion can be visually demonstrated. Figure 4 shows the PEC, curve, and torsional wavefunctions of some CXYZOH molecules. The curves take various values along , and the torsional wavefunctions located in different wells can exhibit different values. This yields large values in the transition across the two wells. This is in contrast to SAMs, where the small displacement from the equilibrium structure contributes to the vibrational wavefunction, and the two states exhibit a similar magnitude. Following Ref. 29, the contribution from the term is small, yielding a linear-like curve with respect to the vibrational coordinate (). The value and vibrational band origin for each torsional state are presented in Table 1. Assuming the state control was possible, the maximum sensitivity of 3.1 Hz would be achieved in CHBrIOH. The PV frequencies from the ground state exceeded more than 1.0 Hz in the highlighted transitions of the CHBrIOH and CFBrIOH molecules. Compared with the CHFClBr molecule with = 1.867 (–3.781) mHz for the C-F stretching (FCCl bending) fundamental modes [71], approximately 100- (50-)fold enhancement was observed in CFClBrOH (186.4 mHz) without including a heavy element. The necessary precisions (the ratio between the PV shift and torsional frequency) for detecting the PV shift are and for CHBrIOH and CFClBrOH, respectively.
These PV enhancements would not be due to the approximation employed in this study. The qualitative accuracy of the reduced dimensional models that consider only large-amplitude motion(s) has been confirmed in previous studies [72, 73, 74, 41, 75]. The PBE0 functional offers an acceptable quality (maximum absolute error of 0.5 kcal/mol from the CCSD(T) level) for the torsional potential energy including halogens (Figure 5 of ref [76]). The torsional coordinates that offer the maximum, zero, and minimum values do not significantly depend on the methodology choice, e.g., the sinusoidal curves of H2X2 (Fig. 1 of Ref. 20), and the torsional curves of dichlorinedioxide and substituted allenes [77]. Although the maximum and minimum values depend on the methodology, the error due to PBE0 would be much smaller than the enhancement, as the maximum difference between at the PBE0 and CCSD levels of theory is approximately 11% for the H2X2 (X = O, S, Se, Te, Po) molecules [70]. The PV enhancement due to the torsional motion was also observed with the CAM-B3LYP* functional [78], as shown in Sec S6 of the SM.
Another advantage of the transitions across the well is that the extremely long lifetime of the torsional excited states is expected. When the torsional wavefunctions are mainly localized to one of the three wells, they may be recognized as exhibiting different vibrational modes from one that is localized to another well. For example, the localized torsional states of #1, #2, and #3 of CClBrIOH can be assigned to hypothetical vibrational modes with different quantum numbers, , and . These nearly forbidden transitions produce a long lifetime of the vibrational excited states, resulting in a very narrow line width for the transition.
60 | 180 | 300 | CHBrIOH | 60 | 180 | 300 | CFBrIOH | ||||
1 | 0 | 0 | 1 | 191 | –259.7 | 1 | 0 | 0 | 1 | 144 | –674.1 |
2 | 0 | 0 | 2 | 376 | –402.7 | 2 | 0 | 1 | 0 | 224 | –682.2 |
3 | 0 | 0 | 3 | 710 | –547.3 | 3 | 0 | 1 | 2 | 260 | –638.1 |
4 | 0 | 0 | 4 | 990 | –746.6 | 4 | 1 | 0 | 0 | 317 | 672.7 |
5 | 0 | 1 | 4 | 1115 | 1556.9 | 5 | 0 | 2 | 2 | 417 | –483.5 |
6 | 1 | 1 | 5 | 1190 | –1277.0 | 6 | 0 | 2 | 3 | 484 | –725.5 |
7 | 1 | 1 | 5 | 1218 | –1629.4 | 7 | - | - | - | 529 | 204.8 |
8 | 1 | 2 | 5 | 1320 | 474.2 | 8 | - | - | - | 629 | –404.7 |
9 | - | - | - | 1358 | –578.0 | 9 | - | - | - | 665 | –172.5 |
60 | 180 | 300 | CFClBrOH | 60 | 180 | 300 | CClBrIOH | ||||
1 | 0 | 0 | 1 | 132 | –82.2 | 1 | 0 | 0 | 1 | 155 | –479.5 |
2 | 1 | 0 | 0 | 151 | 104.2 | 2 | 0 | 1 | 0 | 37 | –168.0 |
3 | 0 | 0 | 2 | 229 | –54.4 | 3 | 1 | 0 | 0 | 133 | 367.0 |
4 | 0 | 1 | 2 | 264 | –33.4 | 4 | 0 | 2 | 2 | 267 | –364.4 |
5 | 2 | 0 | 3 | 360 | 22.4 | 5 | 0 | 2 | 2 | 284 | –319.1 |
6 | - | - | - | 421 | 5.2 | 6 | 2 | 2 | 0 | 364 | 332.7 |
7 | - | - | - | 486 | –21.1 | 7 | - | - | - | 442 | –301.3 |
8 | - | - | - | 584 | –0.8 | 8 | - | - | - | 524 | –234.6 |
9 | - | - | - | 634 | –9.0 | 9 | - | - | - | 559 | –58.1 |
Some torsional transitions appear in the terahertz (THz) range, where lasers are actively developed to achieve stable frequencies [79], as in the case of quantum cascade lasers [80, 81]. However, the transition to the SAM-LAM coupling mode can reach the wavelengths in the infrared range, where CO2 laser can reach. In the case of the H-containing species, PV-sensitive transitions are observed in the infrared range, such as the transition between #1 and #5 of CHBrIOH presented in Table 1. To the best of our knowledge, these substituted methanol molecules have not been synthesized; however, they may be formulated in the gas phase via laser ablation, which is widely employed (e.g., Refs. 82, 83, 84, 85, 86).
V Conclusions
In conclusion, an enhancement mechanism for the PV frequency originating in the torsional motions of substituted chiral methanol molecules was demonstrated. The torsional transition of the CFClBrOH molecule exhibits approximately 100 times more sensitivity to the PV measurement compared with the conventional C-F stretching motion of the CHFClBr molecule. The calculated PV shift of CHBrIOH reached 3.1 Hz. The torsional energy levels were calculated using the exact quantum dynamics method. The PECs and curves of the torsional motion were calculated using a DFT method based on the two-component relativistic eAMF-X2C Hamiltonian. The key idea of the enhancement mechanism is that the torsional states located in different wells can have with the opposite sign.
[…]
Acknowledgements.
The author wishes to acknowledge a financial support from the European Union’s Horizon 2022 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101105452 (QDMAP) and the JSPS KAKENHI (grant no 21K14643). The author used the computer resource offered under the category of General Projects by the Research Institute for Information Technology, Kyushu University, and the FUJITSU Supercomputer PRIMEHPC FX1000 and FUJITSU Server PRIMERGY GX2570 (Wisteria/BDEC-01) at the Information Technology Center, The University of Tokyo. The author thanks Edit Mátyus (Budapest) for valuable discussions and Alberto Martń Santa Daría (Salamanca) for the support in adding the PV module to the TENSINT code. The author thanks Editage (www.editage.com) for English language editing.References
- Miyamoto et al. [2022] Y. Miyamoto, R. Tobaru, Y. Takahashi, A. Hiramoto, K. Iwakuni, S. Kuma, K. Enomoto, and M. Baba, High-resolution spectroscopy of buffer-gas-cooled phthalocyanine, Commun Chem 5, 161 (2022).
- Aiello et al. [2022] R. Aiello, V. Di Sarno, M. G. Delli Santi, M. De Rosa, I. Ricciardi, P. De Natale, L. Santamaria, G. Giusfredi, and P. Maddaloni, Absolute frequency metrology of buffer-gas-cooled molecular spectra at 1 kHz accuracy level, Nat. Commun. 13, 7016 (2022).
- Anderegg et al. [2023] L. Anderegg, N. B. Vilas, C. Hallas, P. Robichaud, A. Jadbabaie, J. M. Doyle, and N. R. Hutzler, Quantum control of trapped polyatomic molecules for eEDM searches, Science 382, 665 (2023).
- Vilas et al. [2024] N. B. Vilas, P. Robichaud, C. Hallas, G. K. Li, L. Anderegg, and J. M. Doyle, An optical tweezer array of ultracold polyatomic molecules, Nature 628, 282 (2024).
- Wu et al. [1957] C. S. Wu, E. Ambler, R. W. Hayward, D. D. Hoppes, and R. P. Hudson, Experimental test of parity conservation in beta decay, Physical Review 105, 1413 (1957).
- Barkov and Zolotorev [1978] L. M. Barkov and M. Zolotorev, Observation of parity nonconservation in atomic transitions, JETP Lett. 27, 357 (1978).
- Barkov and Zolotorev [1979] L. M. Barkov and M. S. Zolotorev, Parity violation in atomic bismuth, Phys. Lett. B 85, 308 (1979).
- Wood et al. [1997] C. S. Wood, S. C. Bennett, D. Cho, B. P. Masterson, J. L. Roberts, C. E. Tanner, and C. E. Wieman, Measurement of parity nonconservation and an anapole moment in cesium, Science 275, 1759 (1997).
- Daussy et al. [1999] C. Daussy, T. Marrel, A. Amy-Klein, C. T. Nguyen, C. J. Bordé, and C. Chardonnet, Limit on the parity nonconserving energy difference between the enantiomers of a chiral molecule by laser spectroscopy, Phys. Rev. Lett. 83, 1554 (1999).
- Ziskind et al. [2002] M. Ziskind, C. Daussy, T. Marrel, and C. Chardonnet, Improved sensitivity in the search for a parity-violating energy difference in the vibrational spectrum of the enantiomers of CHFClBr, Eur. Phys. J. D 20, 219 (2002).
- Berger [2004] R. Berger, Parity-violation effects in molecules, in Theoretical and Computational Chemistry, Vol. 14 (Elsevier, 2004) pp. 188–288.
- Quack et al. [2008] M. Quack, J. Stohner, and M. Willeke, High-resolution spectroscopic studies and theory of parity violation in chiral molecules, Annu. Rev. Phys. Chem. 59, 741 (2008).
- Berger and Stohner [2019] R. Berger and J. Stohner, Parity violation, Wiley Interdiscip. Rev. Comput. Mol. Sci. 9, 1 (2019).
- Bargueño et al. [2008] P. Bargueño, A. Dobado, and I. Gonzalo, Could dark matter or neutrinos discriminate between the enantiomers of a chiral molecule?, EPL 82, 13002 (2008).
- Gaul et al. [2020] K. Gaul, M. G. Kozlov, T. A. Isaev, and R. Berger, Chiral molecules as sensitive probes for direct detection of P-odd cosmic fields, Phys. Rev. Lett. 125, 123004 (2020).
- Zel’dovich [1977] I. B. Zel’dovich, Energy difference between right-hand and left-hand molecules, due to parity nonconservation in weak interactions of electrons with nuclei, Pis’ma Zh. Eksp. Toer. Fiz. 25, 106 (1977).
- Rein et al. [1979] D. W. Rein, R. A. Hegstrom, and P. G. H. Sandars, Parity non-conserving energy difference between mirror image molecules, Phys. Lett. A 71, 499 (1979).
- Hegstrom et al. [1980] R. A. Hegstrom, D. W. Rein, and P. G. H. Sandars, Calculation of the parity nonconserving energy difference between mirror-image molecules, J. Chem. Phys. 73, 2329 (1980).
- Laerdahl and Schwerdtfeger [1999] J. K. Laerdahl and P. Schwerdtfeger, Fully relativistic ab initio calculations of the energies of chiral molecules including parity-violating weak interactions, Phys. Rev. A 60, 4439 (1999).
- Bast et al. [2011] R. Bast, A. Koers, A. S. P. Gomes, M. Iliaš, L. Visscher, P. Schwerdtfeger, and T. Saue, Analysis of parity violation in chiral molecules, Phys. Chem. Chem. Phys. 13, 864 (2011).
- Bouchiat and Bouchiat [1974] Bouchiat and C. Bouchiat, I. parity violation induced by weak neutral currents in atomic physics, Journal De Physique 35, 899 (1974).
- Figgen et al. [2010] D. Figgen, T. Saue, and P. Schwerdtfeger, Relativistic four- and two-component calculations of parity violation effects in chiral tungsten molecules of the form NWXYZ (X, Y, Z=H, F, Cl, Br, or I), J. Chem. Phys. 132, 10.1063/1.3439692 (2010).
- Schwerdtfeger et al. [2003] P. Schwerdtfeger, J. Gierlich, and T. Bollwein, Large parity-violation effects in heavy-metal-containing chiral compounds, Angew. Chem. Int. Ed Engl. 42, 1293 (2003).
- Saleh et al. [2018] N. Saleh, R. Bast, N. Vanthuyne, C. Roussel, T. Saue, B. Darquié, and J. Crassous, An oxorhenium complex bearing a chiral cyclohexane-1-olato-2-thiolato ligand: Synthesis, stereochemistry, and theoretical study of parity violation vibrational frequency shifts, Chirality 30, 147 (2018).
- Sahu et al. [2023] N. Sahu, K. Gaul, A. Wilm, M. Schnell, and R. Berger, Towards detection of molecular parity violation by microwave spectroscopy of CpRe(CH3)(CO)(NO), arXiv: 2303.08263 [physics.chem-ph] (2023).
- Fiechter et al. [2022] M. R. Fiechter, P. A. B. Haase, N. Saleh, P. Soulard, B. Tremblay, R. W. A. Havenith, R. G. E. Timmermans, P. Schwerdtfeger, J. Crassous, B. Darquié, L. F. Pašteka, and A. Borschevsky, Toward detection of the molecular parity violation in chiral ru(acac)3 and os(acac)3, J. Phys. Chem. Lett. 13, 10011 (2022).
- Berger and Stuber [2007] R. Berger and J. L. Stuber, Electroweak interactions in chiral molecules: two-component density functional theory study of vibrational frequency shifts in polyhalomethanes, Mol. Phys. 105, 41 (2007).
- Wormit et al. [2014] M. Wormit, M. Olejniczak, A.-L. Deppenmeier, A. Borschevsky, T. Saue, and P. Schwerdtfeger, Strong enhancement of parity violation effects in chiral uranium compounds, Phys. Chem. Chem. Phys. 16, 17043 (2014).
- Eduardus et al. [2023] Eduardus, Y. Shagam, A. Landau, S. Faraji, P. Schwerdtfeger, A. Borschevsky, and L. F. Pašteka, Large vibrationally induced parity violation effects in CHDBrI+, Chem. Commun. 59, 14579 (2023).
- Lazzeretti and Zanasi [1997] P. Lazzeretti and R. Zanasi, On the calculation of parity-violating energies in hydrogen peroxide and hydrogen disulphide molecules within the random-phase approximation, Chem. Phys. Lett. 279, 349 (1997).
- Bakasov et al. [1998] A. Bakasov, T. Ha, and M. Quack, Ab initio calculation of molecular energies including parity violating interactions, J. Chem. Phys. 109, 7263 (1998).
- Thyssen et al. [2000] J. Thyssen, J. K. Laerdahl, and P. Schwerdtfeger, Fully relativistic coupled cluster treatment for parity-violating energy differences in molecules, Phys. Rev. Lett. 85, 3105 (2000).
- Berger and Van Wüllen [2005] R. Berger and C. Van Wüllen, Density functional calculations of molecular parity-violating effects within the zeroth-order regular approximation, J. Chem. Phys. 122, 10.1063/1.1869467 (2005).
- Berger et al. [2001] R. Berger, M. Gottselig, M. Quack, and M. Willeke, Parity violation dominates the dynamics of chirality in dichlorodisulfane, Angew. Chem. Int. Ed. 40, 4195 (2001).
- Quack and Willeke [2003] M. Quack and M. Willeke, Theory of stereomutation dynamics and parity violation in hydrogen thioperoxide isotopomers 1,2,3HSO1,2,3H, Helv. Chim. Acta 86, 1641 (2003).
- Berger et al. [2003] R. Berger, M. Quack, A. Sieben, and M. Willeke, Parity-Violating potentials for the torsional motion of methanol (CH3OH) and its isotopomers (CD3OH,13CH3OH, CH3OD, CH3OT, CHD2OH, and CHDTOH), Helv. Chim. Acta 86, 4048 (2003).
- Sunaga et al. [2024] A. Sunaga, G. Avila, and E. Mátyus, Variational vibrational states of methanol (12D), J. Chem. Theory Comput. 20, 8100 (2024).
- Meyer and Günthard [1969] R. Meyer and H. H. Günthard, Internal rotation and vibration in CH2=CCl–CH2D, J. Chem. Phys. 50, 353 (1969).
- Bell [1994] S. Bell, Variation of geometry, vibrational frequencies and zero-point energies with internal rotation, J. Mol. Struct. 320, 125 (1994).
- Podolsky [1928] B. Podolsky, Quantum-Mechanically correct form of hamiltonian function for conservative systems, Phys. Rev. 32, 812 (1928).
- Mátyus et al. [2009] E. Mátyus, G. Czakó, and A. G. Császár, Toward black-box-type full- and reduced-dimensional variational (ro)vibrational computations, J. Chem. Phys. 130, 134112 (2009).
- Fábri et al. [2011] C. Fábri, E. Mátyus, and A. G. Császár, Rotating full- and reduced-dimensional quantum chemical models of molecules, J. Chem. Phys. 134, 074105 (2011).
- Moskalev et al. [1976] A. N. Moskalev, R. M. Ryndin, and I. B. Khriplovich, Possible lines of research into weak-interaction effects in atomic physics, Sov. Phys. Usp. 19, 220 (1976).
- Greiner and Muller [1996] W. Greiner and B. Muller, Gauge theory of weak interactions, Vol. 5 (Springer, 1996).
- Barra et al. [1986] A. L. Barra, J. B. Robert, and L. Wiesenfeld, Parity non-conservation and NMR observables. calculation of tl resonance frequency differences in enantiomers, Phys. Lett. A 115, 443 (1986).
- Laubender and Berger [2006] G. Laubender and R. Berger, Electroweak quantum chemistry for nuclear-magnetic-resonance-shielding constants: Impact of electron correlation, PHYSICAL REVIEW A 74, 032105 (2006).
- Weijo et al. [2007] V. Weijo, R. Bast, P. Manninen, T. Saue, and J. Vaara, Methodological aspects in the calculation of parity-violating effects in nuclear magnetic resonance parameters, J. Chem. Phys. 126, 074107 (2007).
- Aucar et al. [2022] I. A. Aucar, Y. Chamorro, and A. Borschevsky, Parity-violating contributions to nuclear spin-rotation interactions and to NMR shielding constants in tetrahedral molecules, Phys. Rev. A 106, 062802 (2022).
- [49] 0.2319 is the default value of the version of the DIRAC code employed in this study. The value of depends on the renormarization prescription, see Ref. 50.
- A Zyla et al. [2020] P. A Zyla, R. M Barnett, J. Beringer, O. Dahl, A. D Dwyer, E. D Groom, C.-J. Lin, K. S Lugovsky, E. Pianori, J. D Robinson, et al., Review of particle physics (rpp 2020), Progress of Theoretical and Experimental Physics 2020, 1 (2020).
- Visscher and Dyall [1997] L. Visscher and K. G. Dyall, Dirac–Fock atomic electronic structure calculations using different nuclear charge distributions, At. Data Nucl. Data Tables 67, 207 (1997).
- Baerends et al. [1990] E. J. Baerends, W. H. E. Schwarz, P. Schwerdtfeger, and J. G. Snijders, Relativistic atomic orbital contractions and expansions: magnitudes and explanations, J. Phys. B At. Mol. Opt. Phys. 23, 3225 (1990).
- Mátyus et al. [2023] E. Mátyus, A. Martín Santa Daría, and G. Avila, Exact quantum dynamics developments for floppy molecular systems and complexes, Chem. Commun. 59, 366 (2023).
- Martín Santa Daría et al. [2021] A. Martín Santa Daría, G. Avila, and E. Matyus, Performance of a black-box-type rovibrational method in comparison with a tailor-made approach: Case study for the methane–water dimer, The Journal of Chemical Physics 154, 224302 (2021).
- Martín Santa Daría et al. [2024] A. Martín Santa Daría, G. Avila, and E. Mátyus, Methane dimer rovibrational states and raman transition moments, Phys. Chem. Chem. Phys. 26, 10254 (2024).
- Frisch et al. [2016] M. J. Frisch, G. W. Trucks, H. B. Schlegel, G. E. Scuseria, M. A. Robb, J. R. Cheeseman, G. Scalmani, V. Barone, G. A. Petersson, H. Nakatsuji, X. Li, M. Caricato, A. V. Marenich, J. Bloino, B. G. Janesko, R. Gomperts, B. Mennucci, H. P. Hratchian, J. V. Ortiz, A. F. Izmaylov, J. L. Sonnenberg, D. Williams-Young, F. Ding, F. Lipparini, F. Egidi, J. Goings, B. Peng, A. Petrone, T. Henderson, D. Ranasinghe, V. G. Zakrzewski, J. Gao, N. Rega, G. Zheng, W. Liang, M. Hada, M. Ehara, K. Toyota, R. Fukuda, J. Hasegawa, M. Ishida, T. Nakajima, Y. Honda, O. Kitao, H. Nakai, T. Vreven, K. Throssell, J. A. Montgomery, Jr., J. E. Peralta, F. Ogliaro, M. J. Bearpark, J. J. Heyd, E. N. Brothers, K. N. Kudin, V. N. Staroverov, T. A. Keith, R. Kobayashi, J. Normand, K. Raghavachari, A. P. Rendell, J. C. Burant, S. S. Iyengar, J. Tomasi, M. Cossi, J. M. Millam, M. Klene, C. Adamo, R. Cammi, J. W. Ochterski, R. L. Martin, K. Morokuma, O. Farkas, J. B. Foresman, and D. J. Fox, Gaussian 16 Revision A.03 (2016), gaussian Inc. Wallingford CT.
- Peterson et al. [2003] K. A. Peterson, D. Figgen, E. Goll, H. Stoll, and M. Dolg, Systematically convergent basis sets with relativistic pseudopotentials. II. small-core pseudopotentials and correlation consistent basis sets for the post-d group 16–18 elements, J. Chem. Phys. 119, 11113 (2003).
- Peterson et al. [2006] K. A. Peterson, B. C. Shepler, D. Figgen, and H. Stoll, On the spectroscopic and thermochemical properties of ClO, BrO, IO, and their anions, J. Phys. Chem. A 110, 13877 (2006).
- Weigend and Ahlrichs [2005] F. Weigend and R. Ahlrichs, Balanced basis sets of split valence, triple zeta valence and quadruple zeta valence quality for H to rn: Design and assessment of accuracy, Phys. Chem. Chem. Phys. 7, 3297 (2005).
- Adamo and Barone [1999] C. Adamo and V. Barone, Toward reliable density functional methods without adjustable parameters: The PBE0 model, J. Chem. Phys. 110, 6158 (1999).
- Ernzerhof and Scuseria [1999] M. Ernzerhof and G. E. Scuseria, Assessment of the Perdew–Burke–Ernzerhof exchange-correlation functional, J. Chem. Phys. 110, 5029 (1999).
- Grimme et al. [2010] S. Grimme, J. Antony, S. Ehrlich, and H. Krieg, A consistent and accurate ab initio parametrization of density functional dispersion correction (DFT-D) for the 94 elements H-pu, J. Chem. Phys. 132, 154104 (2010).
- Saue et al. [2020] T. Saue, R. Bast, A. S. P. Gomes, H. J. A. Jensen, L. Visscher, I. A. Aucar, R. Di Remigio, K. G. Dyall, E. Eliav, E. Fasshauer, et al., The DIRAC code for relativistic molecular calculations, J. Chem. Phys. 152, 204104 (2020).
- [64] DIRAC, a relativistic ab initio electronic structure program, Release DIRAC22 (2022), written by H. J. Aa. Jensen, R. Bast, A. S. P. Gomes, T. Saue and L. Visscher et al. (available at http://dx.doi.org/10.5281/zenodo.6010450, see also http://www.diracprogram.org), accessed Nov. 5, 2022.
- Iliaš and Saue [2007] M. Iliaš and T. Saue, An infinite-order two-component relativistic hamiltonian by a simple one-step transformation, J. Chem. Phys. 126, 64102 (2007).
- Knecht et al. [2022] S. Knecht, M. Repisky, H. J. A. Jensen, and T. Saue, Exact two-component hamiltonians for relativistic quantum chemistry: Two-electron picture-change corrections made simple, J. Chem. Phys. 157, 114106 (2022).
- Dyall [2016] K. G. Dyall, Relativistic double-zeta, triple-zeta, and quadruple-zeta basis sets for the light elements H–Ar, Theor. Chem. Acc. 135, 128 (2016).
- Dyall [2002] K. G. Dyall, Relativistic and nonrelativistic finite nucleus optimized triple-zeta basis sets for the 4p, 5p and 6p elements, Theoretical Chemistry Accounts 108, 335–340 (2002).
- Dyall [2006] K. G. Dyall, Relativistic quadruple-zeta and revised triple-zeta and double-zeta basis sets for the 4p, 5p, and 6p elements, Theoretical Chemistry Accounts 115, 441–447 (2006).
- Sunaga and Saue [2021] A. Sunaga and T. Saue, Towards highly accurate calculations of parity violation in chiral molecules: relativistic coupled-cluster theory including QED-effects, Molecular Physics , e1974592 (2021).
- Rauhut and Schwerdtfeger [2021] G. Rauhut and P. Schwerdtfeger, Parity-violation effects in the vibrational spectra of CHFClBr and CDFClBr, Phys. Rev. A 103, 042819 (2021).
- Senent [2001] M. L. Senent, Ab initio determination of the torsional spectra of acetic acid, Mol. Phys. 99, 1311 (2001).
- Blasco and Lauvergnat [2003] S. Blasco and D. Lauvergnat, Quantum study of the internal rotation of methanol in full dimensionality (1+11d): a harmonic adiabatic approximation, Chem. Phys. Lett. 373, 344 (2003).
- Csaszar et al. [2004] A. G. Csaszar, V. Szalay, and M. L. Senent, Ab initio torsional potential and transition frequencies of acetaldehyde, J. Chem. Phys. 120, 1203 (2004).
- Ragni et al. [2009] M. Ragni, A. Lombardi, P. R. P. Barreto, and A. C. P. Bitencourt, Orthogonal coordinates and hyperquantization algorithm. the NH3 and H3O+ umbrella inversion levels, J. Phys. Chem. A 113, 15355 (2009).
- Tahchieva et al. [2018] D. N. Tahchieva, D. Bakowies, R. Ramakrishnan, and O. A. von Lilienfeld, Torsional potentials of glyoxal, oxalyl halides, and their thiocarbonyl derivatives: Challenges for popular density functional approximations, J. Chem. Theory Comput. 14, 4806 (2018).
- Horný and Quack [2015] L. Horný and M. Quack, Computation of molecular parity violation using the coupled-cluster linear response approach, Mol. Phys. 113, 1768 (2015).
- Thierfelder et al. [2010] C. Thierfelder, G. Rauhut, and P. Schwerdtfeger, Relativistic coupled-cluster study of the parity-violation energy shift of CHFClBr, Phys. Rev. A 81, 032513 (2010).
- Guo et al. [2024] X. Guo, K. Bertling, B. C. Donose, M. Brünig, A. Cernescu, A. A. Govyadinov, and A. D. Rakić, Terahertz nanoscopy: Advances, challenges, and the road ahead, Appl. Phys. Rev. 11, 10.1063/5.0189061 (2024).
- Zeng et al. [2020] Y. Zeng, B. Qiang, and Q. J. Wang, Photonic engineering technology for the development of terahertz quantum cascade lasers, Adv. Opt. Mater. 8, 1900573 (2020).
- Vitiello and Tredicucci [2021] M. S. Vitiello and A. Tredicucci, Physics and technology of terahertz quantum cascade lasers, Adv. Phys. X. 6, 1893809 (2021).
- Tokunaga et al. [2017] S. K. Tokunaga, R. J. Hendricks, M. R. Tarbutt, and B. Darquié, High-resolution mid-infrared spectroscopy of buffer-gas-cooled methyltrioxorhenium molecules, New J. Phys. 19, 053006 (2017).
- Asselin et al. [2017] P. Asselin, Y. Berger, T. R. Huet, L. Margulès, R. Motiyenko, R. J. Hendricks, M. R. Tarbutt, S. K. Tokunaga, and B. Darquié, Characterising molecules for fundamental physics: an accurate spectroscopic model of methyltrioxorhenium derived from new infrared and millimetre-wave measurements, Phys. Chem. Chem. Phys. 19, 4576 (2017).
- Cournol et al. [2019] A. Cournol, M. Manceau, M. Pierens, L. Lecordier, D. B. A. Tran, R. Santagata, B. Argence, A. Goncharov, O. Lopez, M. Abgrall, Y. Le Coq, R. Le Targat, H. Alvarez Martinez, W. K. Lee, D. Xu, P. E. Pottie, R. J. Hendricks, T. E. Wall, J. M. Bieniewska, B. E. Sauer, M. R. Tarbutt, A. Amy-Klein, S. K. Tokunaga, and B. Darquié, A new experiment to test parity symmetry in cold chiral molecules using vibrational spectroscopy, Quantum Electron. 49, 288 (2019).
- Augenbraun et al. [2020] B. L. Augenbraun, J. M. Doyle, T. Zelevinsky, and I. Kozyryev, Molecular asymmetry and optical cycling: Laser cooling asymmetric top molecules, Phys. Rev. X 10, 31022 (2020).
- Persinger et al. [2023] T. D. Persinger, J. Han, A. T. Le, T. C. Steimle, and M. C. Heaven, Electronic spectroscopy and ionization potentials for YbOH and YbOCH3, Phys. Rev. A 107 (2023).
- Dyall et al. [1989] K. Dyall, I. Grant, C. Johnson, F. Parpia, and E. Plummer, GRASP: A general-purpose relativistic atomic structure program, computer physics communications 55, 425 (1989).
- Becke [1993] A. D. Becke, A new mixing of Hartree–Fock and local density-functional theories, The Journal of chemical physics 98, 1372 (1993).
- Stephens et al. [1994] P. J. Stephens, F. J. Devlin, C. F. Chabalowski, and M. J. Frisch, Ab Initio Calculation of Vibrational Absorption and Circular Dichroism Spectra Using Density Functional Force Fields, J. Phys. Chem. 98, 11623 (1994).
- Yanai et al. [2004] T. Yanai, D. P. Tew, and N. C. Handy, A new hybrid exchange–correlation functional using the coulomb-attenuating method (CAM-B3LYP), Chem. Phys. Lett. 393, 51 (2004).
Supporting Information:
Strong parity-violation effects induced by large-amplitude motions:
A quantum-dynamics study of substituted chiral methanols
VI Employed masses
In the variational vibration calculation, the following atomic masses were used for the nuclei; u, u, u, u, u, u, and u.
VII Convergence with respect to the DVR grid points
The convergence with respect to the discrete variable representation (DVR) grid points and functions was tested using the CHFIOH molecule (Table 2). We obtained the convergence 0.1 for vibrational energies and 0.1 mHz for using 37 DVR grid points up to 20th vibrational states. The 37 DVR grid points and functions are employed for the calculations in the main text.
1 | ||||||||
2 | ||||||||
3 | ||||||||
4 | ||||||||
5 | ||||||||
6 | ||||||||
7 | ||||||||
8 | ||||||||
9 | ||||||||
10 | ||||||||
11 | ||||||||
12 | ||||||||
13 | ||||||||
14 | ||||||||
15 | ||||||||
16 | ||||||||
17 | ||||||||
18 | ||||||||
19 | ||||||||
20 |
VIII Tight exponents
The added exponents to the dyall.3zp basis sets are listed in Table 3. These tight exponents are key to the accurate description of the wavefunction in the core region. These exponents were optimized to reproduce the PV matrix elements of the numerical calculation using the GRASP code [87] within the 1% error. A more detailed explanation of the optimization procedure is described in ref [70].
O | F | Cl | Br | I | ||
second tightest | 4.5531279E+01 | 5.8279110E+01 | 7.8714993E+05 | 4.7203082E+02 | 2.4820723E+04 | 1.1662681E+06 |
tightest | 1.9430165E+02 | 2.4929585E+02 | 6.4063687E+06 | 2.1571374E+03 | 1.6203213E+05 | 7.5310209E+06 |
added | 2.8126178E+03 | 5.1263846E+03 | 5.9958092E+07 | 3.6831954E+04 | 1.8253547E+06 | 5.4139599E+07 |
1.9396596E+05 | 1.5303883E+04 | 1.3685631E+06 | 2.8240083E+07 | |||
2.8277597E+04 | 2.4578813E+07 |
IX PECs, wavefunctions, and PV curves
All the target molecules are summarized in Table 4. The configurations (- or -) were determined so that the deepest PEC well is located around .
-configuration | -configuration |
CHFClOH | CFClBrOH |
CHFBrOH | CFClIOH |
CHFIOH | CFBrIOH |
CHClBrOH | CClBrIOH |
CHClIOH | |
CHBrIOH |
Figures 6-10 display the potential energy curves (PECs), 1D torsional wavefunctions, and parity-violating energy () curves of the molecules that are not shown in the main text. In these figures, the blue and green solid lines indicate the odd (ground, 3rd, …) and even-numbered (2nd, 4th, …) torsional state levels. The horizontal axis represents the torsional angle () defined in Eq. 2 of the main text. An example of the direction of the internal rotation of the CXYZ unit is shown in FIG. 3 of the main text. The atomic number increases from atom X to Z, and the wells correspond to , respectively, corresponding to (H-O-C-Y), (H-O-C-Z), and (H-O-C-X), respectively.
Tables 6-10 present the vibrational energies and PV frequencies of the target molecules. In these tables, ( = X, Y, Z = H, F, Cl, Br, I) represents the quantum number that corresponds to the well where the dihedral angle -C-O-H is close to at the corresponding (approximate) values (). “-” refers to mixed states that are delocalized into the three wells.
The CXYZOH molecules can be broadly categorized into the follwoing two types based on their CXYZ units: those with hydrogen (H-containing) and those without hydrogen (H-non-containing). The H-containing molecules may be preferred from the measurement viewpoint because the deep well associated with the hydrogen vibrational frequency, denoted as , allows for high-energy transitions achievable by a CO2 laser (around 1000 ). Contrarily, the H-non-containing molecules benefit from the localized characteristics of their torsional wavefunction, which arise from the proximity of the three local minimum depths. This feature would yield very long lifetimes owing to the predicted extremely narrow linewidth.
60 | 180 | 300 | CHFClOH | ||
1 | 0 | 0 | 1 | 176.5 | 10.84 |
2 | 0 | 0 | 2 | 334.5 | 8.77 |
3 | 0 | 0 | 3 | 628.1 | 6.13 |
4 | 0 | 0 | 4 | 854.3 | 0.88 |
5 | 1 | 0 | 4 | 945.7 | 0.29 |
6 | 1 | 0 | 5 | 1069.6 | 2.24 |
7 | 0 | 1 | 0 | 1116.8 | 2.46 |
8 | 2 | 1 | 5 | 1194.7 | 2.94 |
9 | - | - | - | 1249.5 | 3.40 |
10 | - | - | - | 1380.3 | 2.89 |
60 | 180 | 300 | CHFBrOH | ||
1 | 0 | 0 | 1 | 182.0 | 135.31 |
2 | 0 | 0 | 2 | 345.9 | 113.57 |
3 | 0 | 0 | 3 | 649.3 | 83.71 |
4 | 0 | 0 | 4 | 869.6 | 0.36 |
5 | 1 | 0 | 4 | 936.0 | 28.90 |
6 | 1 | 0 | 5 | 1075.1 | 27.02 |
7 | 2 | 0 | 5 | 1185.4 | 43.66 |
8 | 2 | 1 | 5 | 1244.0 | 27.36 |
9 | - | - | - | 1292.9 | 46.38 |
10 | - | - | - | 1427.5 | 38.58 |
X Torsional quantum number for the CXYZOH molecules
Traditionally, the quantum numbers at the vibrational ground states are defined as zero for all the vibrational modes. However, this convention would be insufficient in reflecting the character of the torsional wavefunction because ‘’ does not offer information on the bottom of the torsion potential at which the wavefunction is localized. Therefore, in this study, we defined the torsional quantum number as the number of the extremums of the vibrational wavefunction, except for the numerical noises, to clarify the character of the wavefunction.
60 | 180 | 300 | CHFIOH | ||
1 | 0 | 0 | 1 | 188.2 | 603.57 |
2 | 0 | 0 | 2 | 357.6 | 522.90 |
3 | 0 | 0 | 3 | 671.3 | 406.44 |
4 | 0 | 0 | 4 | 908.9 | 63.08 |
5 | 1 | 0 | 4 | 963.0 | 154.39 |
6 | 2 | 0 | 4 | 1103.2 | 125.74 |
7 | 2 | 1 | 4 | 1181.1 | 166.44 |
8 | 2 | 1 | 5 | 1229.7 | 172.83 |
9 | - | - | - | 1292.5 | 225.94 |
10 | - | - | - | 1422.9 | 181.43 |
60 | 180 | 300 | CHClBrOH | ||
1 | 0 | 0 | 1 | 185.5 | 87.84 |
2 | 0 | 0 | 2 | 352.7 | 80.38 |
3 | 0 | 0 | 3 | 667.1 | 68.46 |
4 | 0 | 0 | 4 | 931.1 | 47.54 |
5 | 1 | 0 | 4 | 1045.2 | 32.28 |
6 | 1 | 0 | 5 | 1128.2 | 43.21 |
7 | 1 | 1 | 5 | 1192.8 | 28.58 |
8 | 2 | 1 | 5 | 1269.7 | 11.72 |
9 | - | - | - | 1334.0 | 1.29 |
10 | - | - | - | 1458.2 | 0.47 |
60 | 180 | 300 | CHClIOH | ||
1 | 0 | 0 | 1 | 191.6 | 396.41 |
2 | 0 | 0 | 2 | 363.7 | 384.84 |
3 | 0 | 0 | 3 | 686.1 | 359.17 |
4 | 0 | 0 | 4 | 954.3 | 316.16 |
5 | 1 | 0 | 0 | 1073.5 | 305.51 |
6 | 1 | 1 | 4 | 1115.2 | 48.16 |
7 | 1 | 1 | 5 | 1177.1 | 46.05 |
8 | 2 | 1 | 5 | 1275.4 | 96.26 |
9 | 2 | 2 | 5 | 1312.6 | 61.83 |
10 | - | - | - | 1448.7 | 13.87 |
60 | 180 | 300 | CFClIOH | ||
1 | 0 | 0 | 1 | 135.3 | 412.04 |
2 | 1 | 0 | 0 | 167.4 | 352.51 |
3 | 0 | 1 | 0 | 185.6 | 126.32 |
4 | 0 | 1 | 2 | 238.4 | 332.21 |
5 | 2 | 2 | 2 | 365.2 | 126.10 |
6 | - | - | - | 403.9 | 75.32 |
7 | - | - | - | 460.7 | 169.53 |
8 | - | - | - | 571.8 | 110.53 |
9 | - | - | - | 603.5 | 47.60 |
10 | - | - | - | 757.2 | 91.68 |
XI Comparison with CAM-B3LYP* functional
The enhancement mechanism is also examined using another DFT functional optimized for (the contribution only from Br and Cl atoms) in CHFClBr at the CCSD(T) level, which is noted as CAM-B3LYP* [78].
In the CAM-B3LYP approximation [88, 89, 90], electron-electron Coulomb operator into a short-range density functional and a long-range wave-function-based part is expressed by
where is the distance between two electrons and , , , and are the parameters optimized in Ref. 78.
The comparison between the CAM-B3LYP* and PBE0 functionals is shown in Table 11. The maximum deviation is –0.12 mHz (), which is much smaller than the maximum absolute value shown at , –6.86 (–6.77) mHz at the CAM-B3LYP* (PBE0) functionals. The maximum difference in the PEC is relatively large (–57.6 at ), but the torsional angles that show the local minimum and maximum are very similar.
Table 12 lists the PV frequencies and vibrational energies calculated with the CAM-B3LYP* and PBE0 functionals. The relative difference of becomes large when its absolute value of is close to zero (e.g., #4 and #5). However, the differences () are smaller than the most sensitive transition to the PV (between #1 and #5, mHz). In the states in which the vibrational wavefunction is plane-wave-like (i.e., high-energy vibrational states), the values of at the CAM-B3LYP* and PBE0 levels are very close. From this and the comparison shown in Table 11, the difference between CAM-B3LYP* and PBE0 would be mainly due to the vibrational wavefunction. Although the benchmark data is not reported for the CAM-B3LYP*, the validity of the PBE0 function for the torsional PEC is assessed in Ref [76].
(mHz) | PEC () | |||||
CAM-B3LYP* | PBE0 | CAM-B3LYP* | PBE0 | |||
0 | ||||||
9 | ||||||
18 | ||||||
27 | ||||||
36 | ||||||
45 | ||||||
54 | ||||||
63 | ||||||
72 | ||||||
81 | ||||||
90 | ||||||
99 | ||||||
108 | ||||||
117 | ||||||
126 | ||||||
135 | ||||||
144 | ||||||
153 | ||||||
162 | ||||||
171 | ||||||
180 | ||||||
189 | ||||||
198 | ||||||
207 | ||||||
216 | ||||||
225 | ||||||
234 | ||||||
243 | ||||||
252 | ||||||
261 | ||||||
270 | ||||||
279 | ||||||
288 | ||||||
297 | ||||||
306 | ||||||
315 | ||||||
324 | ||||||
333 | ||||||
342 | ||||||
351 | ||||||
360 |
60 | 180 | 300 | () | (mHz) | |||||
F | Cl | H | CAM-B3LYP* | PBE0 | CAM-B3LYP* | PBE0 | |||
1 | 0 | 0 | 1 | ||||||
2 | 0 | 0 | 2 | ||||||
3 | 0 | 0 | 3 | ||||||
4 | 0 | 0 | 4 | ||||||
5 | 1 | 0 | 4 | ||||||
6 | 1 | 0 | 5 | ||||||
7 | 0 | 1 | 0 | ||||||
8 | 2 | 1 | 5 | ||||||
9 | - | - | - | ||||||
10 | - | - | - | ||||||
11 | - | - | - | ||||||
12 | - | - | - | ||||||
13 | - | - | - | ||||||
14 | - | - | - | ||||||
15 | - | - | - | ||||||
16 | - | - | - | ||||||
17 | - | - | - | ||||||
18 | - | - | - | ||||||
19 | - | - | - | ||||||
20 | - | - | - |